# Conformal Convolution and Monte Carlo Meta-learners for Predictive Inference of Individual Treatment Effects

### v4

@@ 3.2 Universal Consistent CCT- and CMC T-Learner
Assuming independent (potential) outcome noise distributions $\mathbf{\boldsymbol{\varepsilon(0)\perp\!\!\!\perp\varepsilon(1)}}$, and that the object space $\mathcal{X}$ is standard Borel, a universal consistent CCT- and CMC T-learner exists (see Algorithm 1 and 2).
@@
The proof for Theorem 1 is given in Appendix A. In short, it uses Theorem 3 of [30], stating that a universal consistent CPS exists and that the convolution of two independent weakly converging sequences weakly convergence to the independent joint distribution.
@@ 4.1 Experimental Setup
Similar to Nie and Wager [5] and Alaa et al. [15], we evaluated our CCT- and CMC T-learners through six different experiments: three using synthetic data (see Section D.1) and three using semi-synthetic data (see Section D.2). The experiments assessed coverage and efficiency of prediction intervals of the CCT-learner and (pseudo) CMC T-learner against WCP (exact, inexact, and naive) [12], CM-Learner (DR) [15], and a CPS Oracle fitted as a standard regression problem on the ITE (see Appendix C). Additionally, we evaluate the probabilistic calibration on six different synthetic dataset settings. All learners utilized the scikit-learn RandomForestRegressor with default hyperparameters. The CMC learners always employed 100 MC samples. Although experiments with other base learners like gradient-boosting machines were conducted, they were omitted as their results were similar to those of random forests. Additional results of the CMC S- and X-learners are included in Appendix E.
@@ 4.2 Results
(a) Probability Integral Transform (PIT)
(b) 90% ITE prediction interval.
Figure 2: Simulation results for setups A and B from Alaa et al. [15].
Coverage and efficiency of prediction intervals
On synthetic datasets, the CCT- and CMC meta-learners consistently demonstrated slightly conservative coverage across various simulations while maintaining high efficiency, indicated by small average interval widths. In contrast, the inexact WCP, while more efficient in some cases, often did so at the cost of significant under coverage. The CM DR learner was consistently more conservative than our proposed methods. Other WCP approaches, with proven conservative validity, were extremely conservative across all confidence levels and experiments, likely due to their over-conservative design, employing either Bonferroni correction or double conformal inference procedure [12].
For the three semi-synthetic datasets, we performed experiments with covariates from real-world data. As here no propensity scores are provided, these resemble a realistic scenario. We estimate the propensity scores using logistic regression. The most extensive experiment was conducted on the 77 settings of the ACIC2016 dataset [31], which contain 58 covariates (3 categorical, 5 binary, 27 counts, 23 continuous) with varying complexity of outcome surfaces, degree of confounding, overlap, and treatment effect heterogeneity. Results, depicted in Figure 4, show that our proposed approaches provide conservatively valid prediction intervals across all settings, with significant conservative coverage in most settings, though less than the CM DR-learner and valid WCP approaches, likely due to violations in the assumption of conditional independence of the noise distributions. Regarding efficiency, the CCT-learner performed best for the reasons previously mentioned. As can be seen in Figure 11, a similar conclusion can be drawn for the other semi-synthetic experiments.
Performance under different relations between noise distributions
@@
We assessed the probabilistic calibration of the predictive distribution by plotting the probability integral transform (PIT) values across different experiments and simulation runs. A probabilistic forecaster is considered calibrated if it produces uniformly distributed PIT values on the test set [31]. Figure 2(a) shows these PIT values for the synthetic experiments from Alaa et al. [15], indicating slight conservative coverage for both the CCT-learner and CMC T-learner, though minimal.
The learners are almost perfectly calibrated for the synthetic datasets from Nie and Wager [5], as shown in Figure 9. In general, the slight conservative coverage could be due to dependent epistemic uncertainty from shared components between outcomes under treatment and control. Setup D was perfectly calibrated with unrelated outcomes under treatment and control (see Figure 9), supporting this hypothesis. Additionally, Figure 10 illustrates the effect of replacing WCPS with CPS, resulting in uncalibrated learners, except for Setup B, which represents a randomized control trial.
@@ 5 Conclusion
In this work, we present the CCT- and CMC meta-learners for estimating the predictive distribution of CATE estimation to reveal information about the underlying ITE, enhancing decision-making confidence. To the best of our knowledge, our frameworks are the first frequentist approaches to offer an ITE predictive distribution. They outperform the state-of-the-art by demonstrating strong experimental coverage, maintaining narrow interval widths and achieving probabilistic calibration on six different datasets, while enabling the generation of predictive non-parametric distributions. We also highlighted how different assumptions about the noise distribution of potential outcomes can significantly impact the coverage of ITE predictive inference.
@@ Acknowledgement
Part of this research was supported through the Flemish Government (AI Research Program). Jarne Verhaeghe is funded by the Research Foundation Flanders (FWO, Ref. 1S59522N).
Appendix A Proof: Existence Universal Consistent CCT- and CMC T-Learner
The following section presents and proves Theorem 1, which states the existence of a universal consistent CCT- and CMC T-learner, i.e., the conformal predictive distribution approaches the true data generating distribution as $n\rightarrow\infty$ under any data-generating distribution. We introduce here Definition 1, taken in verbatim from [30], which lays out our notion of consistency for CPS.
Definition 1 (in verbatim Vovk [30]).
A conformal predictive system $Q$ is consistent for a probability measure $\mathcal{P}$ on Z := X × R if, for any bounded continuous function $f:\mathbb{R}\rightarrow\mathbb{R}$,
[…]
in probability, where:
  • $Q_{n}$ is the predictive distribution $Q_{n}:y\mapsto Q(z_{1},...,z_{n},x_{test},\phi)$ output by Q as its forecast for the label $y_{test}$ corresponding to the test object $x_{test}$ based on the training data $(z_{1},...,z_{n})$, where $z_{i}=(x_{i},y_{i})$ for all $i$.
  • $\mathbb{E}_{\mathcal{P}}(f|X_{test})$ is the conditional expectation of $f(y)$ given $x=x_{test}$ under $(x,y)\sim\mathcal{P}$.
  • $z_{i}=(x_{i},y_{i})\sim\mathcal{P},i=1,...,n$, $z_{test}\sim\mathcal{P}$, and $\phi\sim Uniform(0,1)$, are assumed all independent.
Theorem 1.
A universal consistent CCT- and CMC T-learner exists (see Algorithm 1 and 2). Assuming independent (potential) outcome noise distributions: $\mathbf{\boldsymbol{\varepsilon(0)\perp\!\!\!\perp\varepsilon(1)}}$, and the object space $\mathcal{X}$ is standard Borel.
Proof.
The philosophy behind the MC simulations in the CMC T-learner is to not perform a convolution of both outcome predictive distributions, as this would quickly become an expensive operation that would need to be performed for every example during inference. However, to prove that there exists a universal consistent, we replace this MC sampling in Algorithm 2 with the convolution operation,
[…]
and thus becoming a CCT-learner, which is equivalent to the CMC T-learner approach asymptotically.
As we consider the existence of a universal consistent learner, we no longer need to worry about the covariate shift. We can replace the WCPS with CPS in both the CCT- and CMC T-learner. This is equivalent to setting the propensity score to $0.5$ for every covariate.
By Theorem 3 in [30], we know that there exists universal CPS if the object space $\mathcal{X}$ is standard Borel, meaning that for any probability measure $\mathcal{P}$ on $Z$, the CPS is consistent. Consequently, there exists a universal CPS that can model the true data-generating distribution outcome under treatment $Y^{1}$ and control $Y^{0}$. These distributions are weakly converging distributions. Since we assume that the predictive distributions of the outcome are independent,
$Y^{0}\perp\!\!\!\perp Y^{1}|X,$ (10)
the outcome variables are independent conditionally on the covariate set $X$. Then, by the definition of convolution and the fact that $ITE=Y^{1}-Y^{0}$, the convolution of the true conditional cumulative distribution of outcome under treatment $F_{Y^{1}|X}$ and outcome under negative control $F_{-Y^{0}|X}$ equals the true conditional cumulative ITE distribution,
[…]
Appendix B Conformal Monte Carlo Meta-Learners
B.1 T-learner
Algorithm 2 Fit and inference of CMC T-learner
1:  Assume: $\phi\sim Uniform(0,1)$, $y\in\mathbb{R}$
2:  Input: training covariates $X_{tr}$, labels $y_{tr}$, treatment indicator $W_{tr}$, propensity score function $\pi(x)$, MC samples $n$, and test object $x_{test}$
3:  $X_{tn},X_{cal},y_{tn},y_{cal},W_{tn},W_{cal}=$split$(X_{tr},y_{tr},W_{tr})$
4:  $X_{..,0}$, $y_{..,0}$:= $X_{..}[W_{..}=0]$, $y_{..}[W_{tr}=0]$
5:  $X_{..,1}$, $y_{..,1}$:= $X_{..}[W_{..}=1]$, $y_{..}[W_{tr}=1]$
6:  $\textbf{Q}_{Y^{0}}(x,y,\phi):=$$\textbf{WCPS}(X_{tn,0}$, $y_{tn,0}$, $x$, $\frac{1}{1-\pi(x)}$,$y$,$\phi)$
7:  $\textbf{Q}_{Y^{1}}(x,y,\phi):=$$\textbf{WCPS}(X_{tn,1}$, $y_{tn,1}$, $x$, $\frac{1}{\pi(x)}$, $y$, $\phi)$
[…]
9:  $\textbf{Q}_{ITE}(x_{test},y,\phi)=\textbf{CPS}(repeat(X_{cal}$, $n$  times)))), $\textbf{ITE}_{MC},x_{test},y,\phi)$

### v5

@@ 4 Conformal convolution and conformal Monte Carlo meta-learner
Assuming independent (potential) outcome noise distributions $\mathbf{\boldsymbol{\varepsilon(0)\perp\!\!\!\perp\varepsilon(1)}}$, and that the object space $\mathcal{X}$ is standard Borel, a universal consistent CCT- and CMC T-learner exists (Algorithm 1 and 2).
@@
The proof for Theorem 3 is given in Appendix E.3. In short, it uses Theorem 3 of Vovk (2022), stating that a universal consistent CPS exists and that the convolution of two independent weakly converging sequences weakly convergence to the independent joint distribution.
Limitations of the CCT-learner and CMC meta-learners
The CCT-learner’s finite-sample calibration guarantee hinges on untestable assumptions: (1) known propensity scores (typically estimated in practice); and (2) weighted exchangeability between pseudo ITE conformity scores and the true ITE $ITE_{n+1}$ (unverifiable due to unobserved counterfactuals). While these assumptions are theoretically critical, empirical results across synthetic and semi-synthetic experiments demonstrate near-perfect probabilistic calibration (e.g., dispersion of PIT values in Section 5.2 and PIT plots in Appendix F), suggesting robustness in practice. Further limitations arise from data efficiency trade-offs inherent in split CPS, as well as the sensitivity to base model’s accuracy.
@@ 5.1 Experimental setup
We evaluate our CCT- and CMC meta-learners through 8 different experiments: 6 using synthetic data (Section G.2) and 2 using semi-synthetic data (Section G.3). Both learners utilized the scikit-learn RandomForestRegressor with default hyperparameters. We benchmark our approaches against the state-of-the-art approaches in Table 2, using the hyperparameters as specified in these works. Benchmark approaches, such as DiffPO or GANITE, are occasionally omitted from the evaluation plots as their poor performance messes up the clarity of the results. The CMC meta-learners always employed 100 MC samples. Given the flexibility of the FCCN and NOFLITE approaches to fit complex noise distribution (e.g., on the semi-synthetic EDU experiment), we implement the CCT-learner with both approaches as base-learners.
@@ 5.2 Results
We discuss the performance of the distribution estimations of the ITE and potential outcomes, focusing on the CMC T-learner here. Results for the S- and X-learners can be found in Appendix F as they are similar to the CMC T-learner, together with the empirical results of the prediction intervals.
Figure 2: Simulation results for semi-synthetic experiments EDU and ACIC 2016.
@@
A predictive distribution is considered probabilistically calibrated if it produces uniformly distributed PIT values on the test set. We assessed the probabilistic calibration of the predictive distribution for all experiments by measuring on the test set the dispersion of the PIT values, i.e., the variance of the PIT values, which should be $\frac{1}{12}$ as this is the dispersion of the standard uniform distribution. Besides measuring the dispersion of the predictive distributions, we also inspect the probabilistic calibration visually for the synthetic experiments from Alaa et al. (2023) and the non-normally distributed semi-synthetic EDU experiment, by plotting the observed frequency of PIT values (Figure 9). In general, we observe that potential outcome distributions of CCT- and CMC T-learners, constructed using WCPS, are perfectly probabilistically calibrated given the ground-truth propensity (Figures 3, 4, 5, 6, 7).
In our semi-synthetic experiments, we always estimate our propensity scores using logistic regression to represent a more realistic setting. Results show that we can generate probabilistically calibrated predictive distributions even when the propensity score is estimated by checking the dispersion and distribution of PIT-values (Figure 2, 8, 9).
In general, we see that most related work approaches have difficulty generating probabilistically calibrated distributions (Figure 9). We see that BART and CMGP are reasonably well calibrated when the noise distribution of the experiment is homoscedastic and normally distributed (Figure 9 for Alaa et al. (2023)). However, once the noise becomes heteroscedastic and certainly nonnormal, their calibration deteriorates (Figure 2, 9). We also observe that the CCT-learner can probabilistically calibrate the FCCN and NOFLITE approaches, often generating slightly overconfident predictive distributions.
CRPS performance
[…]
@@ 6 Conclusion
This work presents the CCT- and CMC meta-learners for estimating the predictive distribution of potential outcomes and ITE, enhancing decision-making confidence. To our knowledge, this is the first framework to provide finite-sample guarantees for calibrating predictive distributions in this context. It outperforms related work by demonstrating relatively strong and consistent CRPS scores while achieving probabilistic calibration. Future work could investigate other forms of finite-sample calibration guarantees and the use of our introduced learners in automatic decision-making.
@@ Acknowledgement
Part of this research was supported through the Flemish Government (AI Research Program). Jef Jonkers is funded by the Research Foundation Flanders (FWO, Ref. 1S11525N). Jarne Verhaeghe is funded by the Research Foundation Flanders (FWO, Ref. 1S59522N).
Appendix A Background
A.1 Conformal prediction
Conformal prediction (CP) is a model-agnostic framework that gives an implicit confidence estimate in a prediction by generating prediction sets at a specified significance level $\alpha$ [Vovk et al., 2022]. Notably, this framework allows for (conservatively) valid non-asymptotic confidence predictors under the exchangeability assumption, assuming training/calibration data exchangeability with test data, a slightly less stringent assumption than the IID assumption. The prediction sets in CP are formed by comparing nonconformity scores of examples that quantify how unusual a predicted label is, i.e., these scores measure the disagreement between the prediction and the actual target.
More specifically, we define a prediction interval $\hat{C}_{\alpha}(X_{n+1})$, for test object $X_{n+1}\in\mathcal{X}$, by computing the conformity scores $S_{i}^{y}$, based on conformity score $s$ for each $y\in\mathbb{R}$, using training data $Z_{i}=(X_{i},Y_{i}),i=1,...,n$:
$\begin{split}\hat{C}_{\alpha}&(X_{n+1})=\Bigl{\{}\\ &y\in\mathbb{R}:\frac{|i=1,...,n+1:s((X_{i},Y_{i});Z_{1:n\setminus i}\cup(X_{n+1},y))\geq s((X_{n+1},y);Z_{1:n})|}{n+1}>\alpha\\ \Bigl{\}}&\end{split}$ (6)
This procedure, termed full or transductive CP, is computationally intensive, as it requires refitting of the underlying model for each $y$ in the target domain. Therefore, a more applicable variant of full CP, called inductive or split CP (ICP) [Papadopoulos et al., 2002], offers computational efficiency and enables the integration of CP with ML algorithms, such as neural networks and tree-based models. In ICP, the training sequence $Z_{1:n}=\{(X_{1},Y_{1}),...,(X_{n},Y_{n})\}$ is portioned into a proper training sequence $\{(X_{1},Y_{1}),...,(X_{m},Y_{m})\}$ and a calibration sequence $\{(X_{m+1},Y_{m+1}),...,(X_{n},Y_{n})\}$. We first train a regression model $\hat{y}(x)$ using the proper training sequence. Then, based on nonconformity score $s$, we generate nonconformity scores $S_{i}$ for data points $(X_{i},Y_{i})$ where $m<i\leq n$. The nonconformity score function $s$ leverages the pre-trained regression model, e.g., the absolute error $s=|y-\hat{y}|$. These nonconformity scores together with test object $X_{n+1}$ and significance level $\alpha$ defines:
$\hat{C}_{\alpha}(X_{n+1})=\left\{y\in\mathbb{R}:\frac{|i=m+1,...,n+1|:s(X_{i},Y_{i})\geq s(X_{n+1},y)|}{n+1}>\alpha\right\}$ (7)
Although the conventional ICP yields calibrated prediction intervals, it only offers marginal coverage and constant interval sizes for different covariates. As we want more conditional valid prediction intervals or at least smaller intervals for confident predictions and larger intervals for more "difficult" examples, extensions of ICP are proposed, such as Mondrian CP [Vovk et al., 2003] by adjusting the CP algorithm and conformalized quantile regression (CQR) [Romano et al., 2019] by introducing a clever nonconformity score. Another adaptation of ICP for establishing more adaptive prediction intervals involves employing normalized nonconformity score [Papadopoulos and Haralambous, 2011]. These normalized scores typically use the product of the absolute error $|y-\hat{f}(x)|$ and the reciprocal of a 1D uncertainty estimate $\frac{1}{\hat{u}(x)}$, expressed as $s=\frac{|y-\hat{f}(x)|}{\hat{u}(x)}$, where $\hat{u}(x)$ may be generated by a distinct model which estimates $|y-\hat{f}(x)|$. For a more comprehensive exploration of CP, its extensions and applications, we refer the reader to other works [Vovk et al., 2022, Angelopoulos and Bates, 2022, Fontana et al., 2022, Manokhin, 2022, Angelopoulos et al., 2024].
Appendix B Related work
B.1 CATE meta-learners
Conditional average treatment effect (CATE) estimation is frequently performed using CATE meta-learners. CATE meta-learners are frameworks where base learners are utilized to learn and perform CATE estimation [Künzel et al., 2019]. Some examples are the T-learner, S-learner, X-learner [Künzel et al., 2019], DR-learner [Kennedy, 2022], and R-learner [Nie and Wager, 2021]. We can categorize these learners into two groups: one-step plug-in (indirect) learners and two-step (direct) learners. One-step plug-in learners (S- and T-learners) are approaches that use two regression functions where the difference represents the CATE function and thus is estimated indirectly. In contrast, two-step learners (X-, DR-, and R-learners) output the CATE function directly. Three learners are used in this work: the T-, S-, and X-learner. In the binary treatment case, the T-learner fits two different base learners $\mu_{0}$ and $\mu_{1}$ on data without treatment and with treatment, respectively. The CATE is then estimated as $\hat{\tau}(X)=\hat{\mu}_{1}(X)-\hat{\mu}_{0}(X)$. In contrast, the S-learner fits a single base learner $\mu$ on $[X,W]$ with $W$ the treatment variable. The CATE is then estimated as $\hat{\tau}(X)=\hat{\mu}(X,W=1)-\hat{\mu}(X,W=0)$. The X-learner starts as a T-learner by first fitting $\mu_{0}$ and $\mu_{1}$, however, afterwards a direct estimate of $\tau$ is fitted with $X$ as input and $W(Y-\hat{\mu}_{0}(X))+(1-W)(\hat{\mu}_{1}(X)-Y)$ as target.
B.2 Prediction intervals for ITE
This section provides a more comprehensive overview of the literature on predictive inference of ITE and quantifying related predictive uncertainty through prediction intervals. Uncertainty quantification is a crucial component for treatment effect estimations for reliable decision-making in high-risk environments, such as healthcare [Banerji et al., 2023]. In the past, most proposals for uncertainty quantification in ITE estimation used Bayesian approaches, such as Bayesian additive regression trees (BART) [Hill, 2011] and Gaussian processes [Alaa and van der Schaar, 2017]. The problem with these Bayesian approaches is that they are often uncalibrated without knowing the underlying data-generating process. These Bayesian approaches are also model-specific and are not easily transferred to modern ML approaches.
Recently, some works have tried to resolve this issue by using the CP framework  [Lei and Candès, 2021, Alaa et al., 2023], which can provide non-asymptotic distribution-free coverage guarantees and thus resolves the significant shortcomings of these Bayesian approaches. Lei and Candès [2021] was the first work to propose a conformal inference approach to produce reliable prediction intervals for ITEs. More specifically, they proposed using WCP to account for covariate shifts as the distributions of covariates for treated and untreated subjects can differ from the target population. They proposed three variants of WCP ITE algorithms: (1) Naive WCP, which combines two potential outcome interval estimates, estimated by applying both WCP and a Bonferroni correction to get an ITE prediction interval; (2) Exact nested WCP, which first generates plug-in estimates of ITEs by using WCP and uses these estimates to apply a regular CP; and (3) Inexact nested WCP which replaces the second CP step in Exact WCP with quantile regression which does not provide any coverage guarantees. A significant drawback of the exact and inexact WCP approaches is that they require refitting the model, at least partially, for different confidence levels for the prediction intervals. Additionally, all the proposed approaches based on WCP only provide intervals as uncertainty quantification.
Alaa et al. [2023] propose conformal meta (CM)-learners, which is a model-agnostic framework for issuing predictive intervals for ITEs that applies the inductive CP on top of CATE meta-learners. These CM-learners are applied to the previously discussed direct meta-learners, which use a two-step procedure where nuisance estimators are fitted on the outcome in the first step. In the second step, these nuisance estimators are used to generate a pseudo outcome in which the covariates are regressed on to get point estimates of the CATE. To construct prediction intervals of the ITE, Alaa et al. [2023] propose to use these pseudo outcomes to generate conformity scores together with the estimates of the meta-learner, i.e., CATE estimator. However, their proposed framework requires more stringent assumptions to be valid than the classical ones for estimating CATE. They developed the so-called stochastic ordering framework and showed that CM-learners are marginally valid if their pseudo outcome conformity scores are stochastically larger than the "oracle" conformity scores evaluated on the unobserved ITEs. However, it is difficult to translate this into practice, as the only way to verify this assumption is to observe the true ITEs, which is impossible. Additionally, their method is limited to producing prediction intervals of ITE only. Note that the same stochastic ordering framework could be integrated with our framework to get validity guarantees of ITE intervals under the same assumption as the CM-learners.
B.3 Potential outcome and ITE predictive distributions
Understanding potential outcomes and ITE predictive distributions is crucial for causal inference, enabling a more nuanced estimation of treatment effects beyond point estimates and prediction intervals. Various methodologies have been proposed to estimate these distributions, leveraging probabilistic models, deep learning techniques, and generative approaches. This section explores existing methods for constructing potential outcomes and ITE predictive distributions, highlighting their underlying assumptions, strengths, and limitations. Table 2 presents a comparative summary of these methods.
BART
Bayesian additive regression trees (BART) [Chipman et al., 2010] have been widely used for causal inference, as proposed by Hill [2011]. This approach uses a Bayesian semiparametric modeling procedure to model the posterior distributions of potential outcomes under the assumption of normally distributed noise. Although Hill [2011] does not explicitly model the predictive ITE distribution, it can be obtained by sampling from the counterfactual posterior distribution and assuming independence between potential outcome distributions. This approach for constructing the ITE distribution can be leveraged for any approach that allows sampling of the potential outcome distributions.
CMGP
Alaa and van der Schaar [2017] introduce causal multi-task Gaussian processes (CMGPs) to provide credible intervals with Bayesian coverage. Their approach models potential outcomes as the outputs of a function in a vector-valued reproducing kernel Hilbert space (vvRKHS). A multi-task Gaussian process (GP), with a linear coregionalization kernel as prior over the vvRKHS, is used for learning potential outcome and ITE predictive distributions. They deal with confounding by using a risk-based empirical Bayes method for adapting the multi-task GP prior that jointly minimizes the empirical error in the observed outcome. While Alaa and van der Schaar [2017] does not explicitly model the ITE predictive distribution, the framework can be extended for this purpose, given its reliance on the error function of the normal distribution. However, this extension remains constrained by prior choices and the normality assumption. The ITE predictive distribution in this setting is given by:
$\widehat{ITE}\sim\mathcal{N}(\hat{\tau}(X_{n+1}),V(X_{n+1}))$ (8)
where $\hat{\tau}(x)=(K_{\theta^{*}}^{T}(x)\Lambda_{\theta^{*}}Y_{1:n})^{T}[-1,1]^{T}$ is the CATE estimate, $V(x)=K_{\theta^{*}}(x,x)-K_{\theta^{*}}(x)\Lambda_{\theta^{*}}K_{\theta^{*}}^{T}(x)$ represents the predictive variance, $K_{\theta^{*}}$ the kernel function parameterized by a set of hyperparameters $\theta^{*}$, and $\Lambda_{\theta^{*}}$ represents the inverse covariance matrix of the training data $X_{1:n}$, plus noise (aleatoric uncertainty). Note that while they do not discuss the relation between the potential outcome noise distributions, they assume independence in their experiments. However, we believe their approach allows for encoding different assumptions of the relation by encoding it into a covariance matrix as a prior.
CEVAE
Another approach to estimate potential outcome distributions involves the causal effect variational autoencoder (CEVAE) [Louizos et al., 2017], which employs deep latent variable modeling and variational autoencoders to infer causal effects. While this approach was designed to deal with hidden confounders, it can also be used in settings without hidden confounders; however, consistency issues are noted in the case of misspecified latent variables or complex data distributions [Rissanen and Marttinen, 2021]. This method also assumes a normal predictive distribution for potential outcomes.
GANITE
While designed to generate predictive ITE distributions, the approach of Yoon et al. [2018], termed generative adversarial nets (GAN) for inference of ITE (GANITE), could also generate potential outcome predictive distributions. GANITE first trains a GAN to generate pseudo counterfactual outcomes, which are then used to train a second GAN. This results in a generator $I(x)$ intended to approximate $(Y(0),Y(1))$. A predictive distribution could be derived using kernel density estimation or additional distributional assumptions. However, the joint distribution obtained from this approach is likely unreliable due to GANITE’s inability to capture dependencies between counterfactual noise distributions (Section D) or implicitly assume a joint distribution of the noise.
DKLITE
Deep kernel learning for ITEs (DKLITE) [Zhang et al., 2020] combines deep kernels with posterior regularization to estimate potential outcomes. This method enables the construction of a potential outcome predictive distribution under the normality assumption. While the authors do not discuss the idea of an ITE predictive distribution, this would only be possible by assuming the conditional independence between the counterfactual predictive distributions and applying the convolution operation or approximating it through sampling.
NOFLITE
NOFLITE [Vanderschueren et al., 2023] employs normalizing flows to optimize the likelihood of potential outcome distributions. This method mitigates confounding by transforming covariates into a balanced representation and explicitly estimates the ITE distribution through Monte Carlo sampling of potential outcome distributions. NOFLITE does not make explicit assumptions about noise distributions but implicitly assumes potential outcome noise independence through its sampling procedure. However, the method focuses on optimizing the likelihood of the conditional prior, capturing aleatoric uncertainty while disregarding epistemic uncertainty.
DiffPO
DiffPO [Ma et al., 2024] introduces a causal diffusion model for learning potential outcome distributions. The model incorporates an orthogonal diffusion loss to address confounding. However, its evaluation primarily compares against approaches using MC dropout, which only accounts for approximation-based epistemic uncertainty. While DiffPO effectively quantifies aleatoric uncertainty and evaluates the alignment between estimated and true aleatoric uncertainty, this evaluation may not fully reflect the comparative performance of the model. We see in our evaluations that this approach fails in most settings; however, this could also be due to insufficient hyperparameter tuning.
FCCN
Full adjustment collaborating causal networks (FCCN) [Zhou et al., 2022] estimate conditional potential outcome distributions while addressing covariate shift through latent representation learning. Based on the collaborating networks framework [Zhou et al., 2021], CCN jointly optimizes two neural networks: one for learning cumulative distributions and another for learning their inverses. Under Lipschitz continuity assumptions, the authors prove that the model converges asymptotically to the true outcome distribution.
Appendix C Conformal Monte Carlo meta-learners
We propose two methods for generating MC ITE samples: (1) The MC ITE method independently samples from both conditional potential outcome predictive distributions, subtracting the potential outcome under control from the one under treatment; (2) The pseudo-MC ITE method adapts this by conditioning on the observed treatment value, calculating the ITE sample as the difference between the observed outcome $Y_{i}$ and the conditional potential outcome predictive distribution under control, and vice versa when treatment is not administered. Since the two approaches sample from different covariate distributions, they require a different weighting to be applied to WCPS. While the MC requires the same weighting scheme as in the CCT-learner, the pseudo-MC procedure requires a weighting of $\frac{\pi(x)}{1-\pi(x)}$ and $\frac{1-\pi(x)}{\pi(x)}$ for the counterfactual outcome of the treated and untreated subjects. Algorithm 3 provides a comprehensive overview of this sampling scheme.
Table 3: Formulas to construct MC ITE sample, where W represents the observed treatment action, Y the observed outcome, and $\hat{Y}(0)$, $\hat{Y}(1)$ are the sampled potential outcome from the conditional potential outcome predictive distribution.
$\mathbf{w_{0}(x)}$ $\mathbf{w_{1}(x)}$
MC $\hat{Y}(1)-\hat{Y}(0)$ $\frac{1}{1-\pi(x)}$ $\frac{1}{\pi(x)}$
pseudo-MC $W(Y-\hat{Y}(0))+(1-W)(\hat{Y}(1)-Y)$ $\frac{\pi(x)}{1-\pi(x)}$ $\frac{1-\pi(x)}{\pi(x)}$